Solve equations by factoring
Solve the equation.
500x^4 = -500x^3 + 45000x^2

Answer by ewatrrr(24785)


Hi,
Solving equation by factoring
500x^4 = -500x^3 + 45000x^2
500x^4 + 500x^3 -45000x^2 = 0
500x^2(x^2 + x - 90) = 0
500x^2(x + 10)(x-9) = 0 Note:SUM of the inner product(10x) and the outer product(-9x) = x
500x^2 = 0
x = 0
(x + 10) = 0
x = -10
(x-9) = 0
x = 9